On January 2, 2015, Concord Corporation issued $1,700,000 of 10% bonds at 97 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”)

The bonds are callable at 102 (i.e., at 102% of face amount), and on January 2, 2020, Concord called $1,020,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Concord as a result of retiring the $1,020,000 of bonds in 2020. (Round answer to 0 decimal places, e.g. 38,548.)

1.)

Loss on redemption


2.) Prepare the journal entry to record the redemption.
Answer #1

Discount on bonds 1020000 = 1020000*.03 = 30600

Unamortied discount = 30600*5/10 = 15300

1) Carrying amount of bonds = 1020000-15300 = 1004700

Loss on redemption = (1020000*1.02)-1004700 =35700

2) Journal entry

Date account and explanation Debit Credit
Bonds payable 1020000
Loss on redemption 35700
Discount on bonds payable 15300
Cash 1040400
(To record redemption)